Brief Exercise 18-11

For Ivanhoe Company, actual sales are $1,310,000, and break-even sales are $1,061,100.

Compute the margin of safety in dollars.
Margin of safety $

Compute the margin of safety ratio.

Margin of safety ratio

%
Answer #1

Solution

Ivanhoe Company

Computation of margin of safety in dollars:

Margin of safety in dollars = actual sales – break-even sales

Actual sales = $1,310,000

Break-even sales = $1,061,100

Margin of safety = 1,310,000 – 1,061,100 = $248,900

Margin of safety ratio = (margin of safety sales/actual sales) x 100

= (248,900/1,310,000) x 100 = 19%